zarkam21:

Calculate the volume (in cm3) of 66 g of water at 1 ∘C.

zarkam21:

zarkam21:

@Vocaloid

Vocaloid:

oh lordy anyway since density = mass/volume then 1000kg/m^3 = (66/1000)kg / V solve for V

zarkam21:

1000kg/m^3 = (66/1000)kg / V 66?

Vocaloid:

hm not quite since 1000 = (66/1000)/V we can swap the V and the 1000 to get V = (66/1000)/1000 of course, that gives you cubic meters so you'd then have to multiply by 10^6 to get it into cm^3

Vocaloid:

like, as a general rule if you have A = B/C you can swap A and C and the expression will remain equal

zarkam21:

6.6 x 10^-5

Vocaloid:

hm. (66/1000)/1000 gives you the units in m^3 multiplying it by 10^6 puts it in cm^3 so actually when you do the math it does end up being 66 whoops making your original sol'n correct

Vocaloid:

which actually does make sense, as a general rule 1g of water is about 1 cm^3 under reasonable conditions

zarkam21:

Calculate the volume (in cm3) of 66 g of water at −1 ∘C. FOr this it would be the same right just with a -1 degree celsius

Vocaloid:

oh actually this is very different since it's below the freezing point and water becomes much less dense looking @ the graph the density at 1 deg. celcius is like 918 kg/m^3 (I remember we had some trouble w/ that last time >>)

Vocaloid:

and repeating the calculations 918 kg/m^3 = (66/1000)kg / x then multiply x by 10^6 to get it in cm^3

zarkam21:

60588000

zarkam21:

so 6.0588 x 10^3 :S

zarkam21:

6.06? if rounded

Vocaloid:

hm not quite solving 918 kg/m^3 = (66/1000)kg / x for x gives us about 0.000071895 m^3 then that times 10^6 is about 72g

Vocaloid:

* 72 cm^3 not g

Vocaloid:

again that's assuming I read the density correctly (this graph is weird)

zarkam21:

wait 65/1000 = .066 and then that times 918 , i got 60.588

zarkam21:

what did I do wrong

Vocaloid:

hm. to isolate 918 kg/m^3 = (66/1000)kg / x you wouldn't multiply 918 and (66/1000) together you would actually swap x and 918 to get x = (66/1000)/918

zarkam21:

oh okay I see

Vocaloid:



zarkam21:

What is the percent change in volume when comparing the volume of 66 g of water at 1 ∘C and at −1 ∘C?

zarkam21:

okay so 72 and 66

zarkam21:

72-66=6 out of 100

zarkam21:

.06 = 6%

Vocaloid:

hm this question isn't worded well but let's assume the 66 is the starting point and the 72 is the ending point then it becomes (72-66)/66, then * 100 to convert to a percentage (since we are comparing from the baseline 66 we divide by 66 not 100)

Vocaloid:

percent change is generally (final - initial)/(initial) * 100

zarkam21:

9.1

zarkam21:

9%

Vocaloid:

yeah good (I think they give you 1 sig fig with the temperature so 9%?)

zarkam21:

it showed as wrong : Use the same equation from Part A, and use the volume at 1 ∘C as the initial value, and the volume at −1 ∘C as the final value. Round to one significant digit.

Vocaloid:

oh it wants the other way around that's cool

Vocaloid:

(66-72)/72 * 100 then

zarkam21:

-8?

Vocaloid:

hoo boy here's the tricky part b/c some formulas consider percent change to be always positive, some don't

Vocaloid:

if it shows you the formula, is there an absolute value sign in the numerator?

zarkam21:

Vocaloid:

ok cool just checking, there's no absolute value so try -8 percent

zarkam21:

Since the density of ice is less than the density of liquid water, the volume of the same mass of water must be more. You are told to indicate the change when water freezes, so the value should be positive.

zarkam21:

Gave me this :S

zarkam21:

so just 8 ?

Vocaloid:

wait what then 9 should have been right the first time :S

zarkam21:

Its not idk

zarkam21:

:/

Vocaloid:

unless the density we used was wrong rendering this whole calculation meaningless :S

zarkam21:

Ugh needs to be submitted by 11:59 and I have two more relative questions to this :S

zarkam21:

Antarctica contains 26.8 million cubic kilometers (km3) of ice. Assume that the average temperature of this ice during the summer months is -20 ∘C. If all of this were heated to 4 ∘C and melted to form water, what volume of liquid water would form?

zarkam21:

We can move to this one. Time is running out so

zarkam21:

its useless sticking to one questioln , you know

zarkam21:

I got it right. It was 8 :S

zarkam21:

d=m/v in this right

Vocaloid:

uh yeah

Vocaloid:

we kind of have to estimate here -20 degrees Celcius the density is about 920kg/m^3 4 degrees celcius the density is about 1000kg/m^3

Vocaloid:

so uh, if we have 26.8 km^3 of ice then that's 920 * 26.8 = 24656 kg of ice then converting that back to volume w/ the new density 24656 kg/x = 1000 x = 24.7 m^3 my best attempt

zarkam21:

The volume of Antarctic ice is given as millions of cubic kilometers rather than just cubic kilometers.

Vocaloid:

oh *** I forgot

zarkam21:

so use mill prob

Vocaloid:

well shouldn't the # be the same either way? :S 24.7 km^3?

zarkam21:

idk The volume of Antarctic ice is given as millions of cubic kilometers rather than just cubic kilometers.

zarkam21:

thats what it says

Vocaloid:

26.8 km^3 * 10^9 * 920 gives 24656000000000kg 24656000000000kg/x = 1000 kg/(m^3) gives 24656000000 m^3 which is 24.7 km^3

zarkam21:

A 1.00-L sample of water is heated from 1 ∘C to 100 ∘C. What is the volume of the water after it is heated?

Vocaloid:

density at 1 celc. is like 1000kg/m^3 density at 100 celc is like 958ish kg/m^3

zarkam21:

lets see if we can get this done in 4 mins lol

zarkam21:

I got the other one i tried 24.7 * 10^7 and that was it

Vocaloid:

therefore: 1 L = 0.001 cubic meters so 1L = 0.001 * 1000 = 1 kg at 100 deg celc 1 kg / x = 958kg/m^3 x = 0.0010438 m^3 ???

zarkam21:

what would that be in sig figs "S

Vocaloid:

they want 3 so 0.00104??

zarkam21:

0.01?

Vocaloid:

0.00104 would have three sig figs

zarkam21:

thats wrong ughh

Vocaloid:

oh *** they want it in L p robably

Vocaloid:

1.04?

zarkam21:

Nvm I got it. Thank you sooo much . I really appreciate it

Vocaloid:

whew ;;